w^2-1/2w-6÷w^2+w-2/3w-9
1
Expert's answer
2021-10-22T00:44:54-0400

"w^2-1\/2w-6\u00f7w^2+w-2\/3w-9"


"w^2-1\/2w-6\/w^2+w-2\/3w-9"


Simplify


"w^2-\frac{1}{2}w-\frac{6}{w^2}+w-\frac{2}{3}w-9"


"w^2-\frac{1}{6}w-\frac{6}{w^2}-9"


"Lcm =2\u00d7w^2\u00d73=6w^2"


"\frac{6w^4-w^3-36-54w^2}{6w^2}"


Hence = "\frac{6w^4-w^3-54w^2-36}{6w^2}"